# Rapid cargo verification with cosmic ray muon scattering and absorption tomography

### v1

@@ 6 Conclusions
In this work, we used Monte Carlo simulations based on the GEANT4 package to investigate the ability of cosmic ray muon tomography to perform prompt characterization of shipping container contents and verification of customs declarations. We developed a method of combined scattering and scattering-to-stopping ratio data analysis to improve the material discrimination ability of cosmic ray muon tomography. This method was applied to analyze simulated datasets for several cargo materials. We demonstrate how integrating scattering and scattering-to-stopping ratio data into a two-dimensional representation enables clearer differentiation between material classifications. The developed approach leverages the combined information from both types of data to enhance the accuracy of material discrimination in muon tomography applications, providing a powerful tool for detecting smuggled goods disguised as other items. For the specific case of tobacco smuggling, the combined analysis of scatter-absorption data allows for the accurate discrimination between tobacco and paper towel rolls with 5.5 accuracy for detector spatial resolution (FWHM) of 0.235 mm, 4.5 for 1.175 mm resolution (FWHM), and 3.5 which corresponds to 99.95% discrimination accuracy for 2.35 mm spatial resolution (FWHM) within a 10-second scanning time.
The ability to rapidly characterize cargo materials comes from the analysis of muon scattering and absorption data for large amounts of cargo material weighing several tons, resulting in high statistical precision. Thanks to the muon penetration power and the developed method of scatter-absorption combined analysis, cosmic-ray muon tomography can be an effective solution for non-invasive cargo inspections, aiding in the fight against contraband trafficking.
@@ 7 Acknowledgments
This work was partially funded by the EU Horizon 2020 Research and Innovation Programme under grant agreement no. 101021812 (“SilentBorder”).

### v3

@@ 6 Conclusions
In this work, we used Monte Carlo simulations based on the GEANT4 package to investigate the ability of cosmic ray muon tomography to perform prompt characterization of shipping container contents and verification of customs declarations. We developed a method of combined muon scattering and absorption data analysis to improve the material discrimination ability of cosmic ray muon tomography. This method was applied to analyze simulated datasets for several cargo materials. We demonstrate how integrating scattering and absorption data into a two-dimensional representation enables clearer differentiation between material. The developed approach leverages the combined information from both types of data to enhance the accuracy of material discrimination in muon tomography applications, providing a powerful tool for detecting smuggled goods disguised as other items. For the specific case of tobacco smuggling, the combined analysis of scatter-absorption data allows for the accurate discrimination between tobacco and paper towel rolls with 5.5 accuracy for detector spatial resolution (FWHM) of 0.235 mm, 4.5 for 1.175 mm resolution (FWHM), and 3.9 which corresponds to 99.95% discrimination accuracy for 2.35 mm spatial resolution (FWHM) within a 10-second scanning time.
The ability to rapidly characterize cargo materials comes from the analysis of muon scattering and absorption data for large amounts of cargo materials weighing several tons, resulting in high statistical precision. Thanks to the muon penetration power and the developed method of scatter-absorption combined analysis, cosmic-ray muon tomography can be an effective solution for non-invasive cargo inspections, aiding in the fight against contraband trafficking.
@@ 7 Acknowledgments
This work was funded by the EU Horizon 2020 Research and Innovation Programme under grant agreement no. 101021812 (“SilentBorder”).